# Harmonia: A Multi-Agent Reinforcement Learning Approach to Data Placement and Migration in Hybrid Storage Systems

### v1

@@ 4.3. Coordination Between Harmonia’s Agents
We achieve implicit coordination between the data-placement and data-migration agents through (1) their rewards, and (2) the effects of their actions on the HSS environment. The data-placement agent’s reward (See Equation 2) encourages effective utilization of the fast device, which is possible only if the data-migration agent proactively migrates cold data to the slow device and hot data to the fast device. The data-migration agent receives a delayed reward (See Equation 3) that is based on access latencies of $n$ incoming I/O requests after moving several migration candidates to their target devices. In this way, each agent’s decisions influence the policy learned by the other agent (See second key challenge in §4.1).
Mitigating Conflicting Decisions. To prevent conflicting placement and migration decisions, which can lead to ping-pong movement of pages across storage devices, we use two key techniques. First, the data-migration agent prioritizes the monitoring of pages that have high access and migration intervals, and ignores recently placed or migrated pages (e.g., if the page’s access interval is 1, i.e., it was recently placed). Second, we add a small penalty based on access and migration intervals to the migration agent’s reward (See Equation 3) to discourage frequent page movements. These mitigation techniques prevent the agents from making conflicting decisions.
@@ 4.4. Harmonia: Design
Data-Placement Agent  1. An RL agent that places I/O request data in the best-fit storage device in the HSS 2 by considering I/O request characteristics and current HSS conditions. The data-placement agent consists of two separate threads. First, the data-placement training thread  3 is a background process in which the agent is trained. Second, the data-placement inference thread  4 is the only foreground thread in Harmonia. In this main decision thread, the data-placement agent decides the target storage device for incoming I/O request data in the critical path.
Data-Migration Agent 5. We implement the data migration using two background threads to avoid interference with the I/O request handling in the critical path. First, the data-migration training thread  6 is used for training the agent. Second, in the data-migration inference thread  7, the agent performs three key tasks: (i) continuously monitors previously-placed pages in the HSS. (ii) determines migration candidates and their target devices. (iii) pushes the metadata of the migration candidates into the migration queue  8. A page is not a migration candidate if the target device determined by the migration agent is the same as its current device.
Migration Queue 8. A queue that stores the metadata of the migration candidates identified by the data-migration agent. Each migration queue entry stores the logical page address and the target device identifier (e.g., 0 for slow device) of each page. We configure the migration queue length to migrate the pages in a timely manner during system idle times and avoid stale migration candidates. Based on empirical analysis (see §6.1), we set the migration queue size to 10 in our experiments.
Migration Thread 9. A background process that continuously monitors the migration queue and migrates pages across devices in the HSS during system idle times. For high-intensity workloads where the system idle times can be very short, the migration thread checks if pages related to incoming read requests are already present in the migration queue. For these pages, it issues low-priority writes (nvme, ; sata, ) to immediately schedule them for migration.
@@ 4.5. Design of Harmonia’s RL Agents
Inference Thread. Harmonia’s data-placement and data-migration agents use the inference thread to make decisions while collecting experiences to use as training data. In the inference thread, the RL agent (1) observes the input features (state) a from the I/O request data and HSS, (2) uses its inference network b to determine the action c for the I/O request, (3) selects the action d that maximizes its long-term rewards, and (4) collects the state, action and reward e as experiences f in an experience buffer g.
Training Thread. In this background thread, Harmonia uses collected experiences h from the experience buffer to train the training network i and optimize the data-placement or data-migration policy. The training thread periodically (e.g., every 1000 requests) transfers the weights of the training network to the inference network for the inference network to adapt to the current workload and HSS conditions.
@@
Hyper-Parameter Tuning. We improve Harmonia’s accuracy by tuning its hyper-parameters (paine2020hyperparameter, ) using cross-validation (arlot2010survey, ) of different values. Harmonia’s agents need only one-time offline hyper-parameter tuning. Table 3 shows the hyper-parameter values chosen based on empirical analysis.
@@
Exploration vs. Exploitation. For every evaluated workload, Harmonia’s data-placement and data-migration agents start with no prior knowledge and make random initial decisions (explore), and use their experiences (exploit) to gradually make optimal placement and migration decisions. To balance exploration and exploitation, we use the $\epsilon$-greedy policy (tokic2011value, ): the predicted best action is selected with (1-$\epsilon$) probability, and a random action with $\epsilon$ probability.
Convergence of RL Agents in Harmonia. Figs 8(a) and  8(b) show the convergence of Harmonia’s data-placement and data-migration policies. We show the convergence in terms of reduction in training loss for the first 10000 I/O requests in YCSB-B (see Table 5) on a performance-optimized HSS.
Figure 8. Convergence of Harmonia’s (a) Data-Placement and (b) Data-Migration RL agents. Convergence is shown in terms of reduction in training loss for each RL agent. We show the first 10000 I/O requests in this figure.
We make two key observations. First, the data-placement policy converges in less than 6000 I/O requests for this workload without any prior training. Second, the data-migration policy takes longer to converge (around 9000 I/O requests) due to the complexity of the state space, and indirect and delayed reward to the data-migration agent. We conclude that the convergence of Harmonia’s agents indicates that they learn to coordinate their actions and make optimal decisions.
@@ 4.6. Overhead Analysis
Harmonia has low inference and training latencies because of its (1) light-weight RL agent design, (2) separation of training and inference, and (3) background execution of multiple threads (e.g., data-migration inference thread). Harmonia’s two RL agents each have identical training and inference networks. We perform training and inference on host CPU because the networks are light-weight, and the neural network weights fit in the on-chip caches of the CPU in our evaluated system.
@@
The storage overhead in Harmonia constitutes (1) the neural networks in the RL agents, (2) the experience buffer, (3) the migration queue, (4) a delayed reward for the migration agent, and (5) address mapping information. First, the weights of the training and inference networks are represented using a half-precision floating-point format. With 90 16-bit weights, each neural network requires $\sim$1.5 KiB of memory. We use the same network architecture for the four networks (training and inference networks for two agents), which takes a negligible 6 KiB of memory. Due to the neural networks’ negligible size, we store their weights in the CPU caches. Second, we use an experience buffer to train each RL agent. Each experience buffer is located in the DRAM, and consumes 100 KiB to store 1000 experiences. Hence, the total storage overhead of two experience buffers is 200 KiB. Overall, Harmonia’s RL agents require 206 KiB of memory. Third, the migration queue holds the logical block address (32 bits) and target device ID (4 bits) for each queue entry. We set the migration queue size to 10 in our experiments. In total, we need 50 bytes of DRAM for the migration queue. Fourth, we store the latencies of 50 incoming I/O requests temporarily to assign a delayed reward to the migration agent, which consumes  100B. Fifth, we maintain the address mapping information for the underlying storage devices (tsukada2021metadata, ) in the HSS. Harmonia requires 32 bits to store the state features for each page (see Table 2). This overhead is less than $\sim$0.8of the total storage capacity when using a 4-KiB data placement granularity.
@@ 5. Evaluation Methodology
We evaluate Harmonia on a real system with an HSS with two (dual-HSS), three (tri-HSS), and four (quad-HSS) devices. The dual-HSS configurations include: (1) performance-optimized HSS, which combines a high-end SSD (inteloptane, ) with a mid-range SSD (intels4510, ), and (2) cost-optimized HSS, which pairs the high-end SSD (inteloptane, ) with a low-end HDD (seagate, ). We implement a custom block device driver in Linux for Harmonia to interface with the underlying storage devices in the HSS. We implement Harmonia using the TF-Agents (TFAgents, ) library. We plan to open-source Harmonia’s implementation to aid future research on HSS. Table 4 provides the details of our real system and the storage devices in the HSS configurations.
@@ 1 TB, SATA 6Gb/s 7200 RPM
high-end (H) & middle-end (M)
& low-end SSD (LSSD)
& low-end HDD (L)
Baselines. We compare Harmonia against four state-of-the-art heuristic- and machine-learning-based data-placement and migration techniques, (1) CDE (matsui2017design, ), (2) RNN-HSS (adapted from  (doudali2019kleio, )), (3) K-SVM (shetti2019machine, ), and (4) Sibyl (singh2022sibyl, ). We evaluate four extended approaches, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM to explore the design space of holistic data-management techniques. We compare Harmonia and prior approaches with two ideal approaches: (1) Fast-Only, and (2) Oracle (meswani2015heterogeneous, ) (See § 3 for details). The Oracle serves as a reference to evaluate the accuracy of Harmonia’s decisions as it makes optimal decisions based on knowledge of future access patterns.
Workloads. We select seventeen data-intensive storage workloads from SYSTOR ’17 (lee2017understanding, ), RocksDB traces (yadgar2021ssd, ), Yahoo! Cloud Serving Benchmark (YCSB) suite (cooper2010benchmarking, ), and MLPerf Storage (mlperfstoragebenchmarks, ) that are collected from real enterprise and datacenter environments. The average workload size is approximately 50000x the fast device capacity. These workloads are chosen to represent diverse I/O access patterns with different read and write ratios, I/O request sizes, and inter-request arrival times. Table 5 reports the characteristics of the chosen workloads. In our evaluation, each workload runs in a separate thread.
@@
Avg. Inter-request
Arrival Time ($\mu$s)
@@
To evaluate Harmonia under real-world scenarios, we generate six multi-programmed workloads by running multiple workloads concurrently (see Table 6 for constituent workloads and their characteristics). We choose these multi-programmed workloads based on three key factors: (1) combination of read-intensive, write-intensive and mixed workloads, (2) number of concurrent workloads required to increase the intensity (i.e., shorter inter-request arrival times) of I/O requests, and (3) combination of different request sizes (See Table 6 for workload characteristics). Each constituent workload is executed using a single thread. For example, mix6 runs on eight concurrent threads for the eight constituent workloads. Multi-programmed workloads typically exhibit higher I/O request intensity and varied access patterns.
@@ 6.1. Performance Analysis
Average I/O Request Latency. Figs. 9(a) and  9(b) show the performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM, Harmonia and Oracle on performance-optimized and cost-optimized HSS configurations. Performance is shown in terms of average request latency normalized to Fast-Only.
Figure 9. Performance of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS configurations. Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make four key observations. First, Harmonia consistently outperforms all the prior and extended techniques in both HSS configurations. In performance-optimized (cost-optimized) HSS, Harmonia provides an average performance improvement of 46.4% (46.7%), 47.9% (57.7%), 43.3% (51.8%), 49.5% (31.7%), 47.6% (63.7%), 49.5% (60.0%), 36.7% (51.4%) and 45.0% (53.8%) over CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM respectively.
Second, Harmonia bridges the performance gap between the best-performing prior technique, Sibyl, and Oracle by 64.2% (64.3%) on average in performance-optimized (cost-optimized) HSS.
Third, Harmonia shows significant performance benefits over prior approaches in read-intensive workloads (e.g., RocksDB, MLPerf) compared to write-intensive workloads. For example, Harmonia outperforms Sibyl by 57.5% (43.2%) in read-intensive workloads and 41.0% (39.9%) in write-intensive workloads. In read-intensive workloads, unlike Sibyl, Harmonia proactively migrates frequently-read pages to the fast device to improve read performance even when there are no writes from the application. In write-intensive workloads, while both Harmonia and Sibyl use RL-based data-placement policies, Harmonia’s data-migration agent, in coordination with the data-placement agent, further improves performance by frequently migrating cold data from the fast device to accommodate new incoming requests.
Fourth, Harmonia shows higher performance gains over prior approaches (other than Sibyl) in cost-optimized HSS than in performance-optimized HSS. This is because (1) data migration in cost-optimized HSS is more expensive than in performance-optimized HSS due to the large latency gap between the two devices, (2) prior heuristic and supervised learning approaches perform data migration during I/O request handling, which adds significant latency overheads, and (3) Harmonia avoids data eviction by freeing up space in the fast device with proactive migrations and performing migrations during system idle times. Sibyl’s RL agent places only performance-critical data in the fast device, reducing the number of data evictions in the critical path.
End-to-End Throughput. Figs. 10(a) and 10(b) show the end-to-end throughput of Harmonia and other evaluated approaches on performance-optimized and cost-optimized HSS, respectively. Throughput is shown in IOPS normalized to that of Fast-Only.
@@
We make three key observations. First, Harmonia improves average end-to-end throughput by 87.8% (210.2%), 101.3% (298.2%), 98.4% (213.1%), 49.4% (156.2%), 91.7% (362.3%), 99.8% (335.5%), 61.7% (160.6%) and 87.8% (312.6%) over CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM in performance-optimized (cost-optimized) HSS. Second, Harmonia achieves an average of 71.7% (62.2%) of the throughput of Oracle on performance-optimized (cost-optimized) HSS. Third, similar to our observation from Figs.  9(a) and 9(b), Harmonia’s throughput gain over Sibyl is higher in read-intensive workloads (i.e., RocksDB) than in write-intensive. This is because, unlike Sibyl, Harmonia proactively migrates data even when there are few updates from the application, resulting in lower latencies and higher throughput.
Tail Latency. Figs. 11(a) and 11(b) show the 99th and 99.99th percentile of I/O latencies (tail latency) of Harmonia and baselines on performance-optimized and cost-optimized HSS, respectively.
@@
We make three key observations. First, Harmonia significantly reduces tail latency compared to all prior approaches across the three workloads. Harmonia reduces the 99th (99.99th) percentile I/O latency by 32.3% (25.3%), 48.4% (45.3%), 49.3% (46.2%), 61.8% (58.5%), 77.7% (75.6%), 92.2% (89.3%), 95.1% (92.7%) and 96.7% (97.5%) on average over Sibyl, CDE, CDE+RL-Migr, SAPM, Sibyl+K-SVM, Sibyl +RNN-HSS, K-SVM and RNN-HSS on performance-optimized (cost-optimized) HSS. Harmonia has lower tail latency because, unlike prior techniques, Harmonia performs data migration during system idle times without interfering with ongoing I/O requests. Second, only ideal policies (i.e., Oracle and Fast-Only) policies show lower tail latency than Harmonia. Third, K-SVM, RNN-HSS, Sibyl+K-SVM and Sibyl+RNN-HSS show very high tail latencies because they perform a large volume of migration at fixed intervals, which adds a significant latency overhead to the ongoing I/O requests.
@@
Performance Sensitivity to Migration Queue Size. Figure 12(a) shows the impact of migration queue size on HSS performance. Performance is shown in terms of average I/O latency normalized to Fast-Only.
We make three key observations. First, the optimal migration queue size is 10 for our evaluated workloads. This may vary slightly across different workloads and HSS configurations. Second, low queue depths (i.e., less than 5) may result in many migration candidates being dropped, which leads to performance degradation. If the migration queue size is 0, Harmonia shows poor performance because it does not perform any data migration. Third, larger-sized queues can hold more migration candidates, but the oldest candidates may become stale (i.e., past migration decisions may not correlate with current access patterns and HSS conditions), which adversely impacts HSS performance.
Performance Sensitivity to Migration Agent’s Reward. Figure 12(b) shows the performance impact of the number of I/O requests considered for the migration agent’s delayed reward (See Equation 3). Performance is shown in terms of average I/O request latency normalized to Fast-Only policy.
@@
We make three key observations on the optimal number of requests for the delayed reward. First, the optimal number of requests is 50 in our evaluated configuration, but may vary slightly across different workloads and HSS configurations, and depends on many factors including (1) inter-request times of incoming I/O requests after the data migration, (2) workload access patterns (e.g., access frequency and reuse distance of a page), and (3) current system conditions (e.g., device capacity utilization) Second, considering fewer requests (e.g., less than 10) may not provide the right reward to the agent because the effect of data migrations is not seen immediately in the placement latencies of these few requests. Third, if we consider more I/O requests (e.g., 250) for the reward, we see lower performance because of two reasons: (i) the impact of data migration is diluted if we consider a large number of incoming requests, and (ii) I/O requests that arrive too far into the future after data migration may not see a direct impact.
Performance on HSS with Persistent Memory (PMEM) Devices. To demonstrate the Harmonia’s extensibility to emerging persistent memory technologies, we evaluate the performance of Harmonia and baselines on an HSS with a persistent memory device (intelpmem200, ) and a high-end SSD (inteloptane, ) (see Table 4 for device characteristics). We emulate the persistent memory device using RAMDisk because these emerging devices are not easily available.
Fig. 13 shows the performance of Harmonia and baselines on an HSS with PMEM device and high-end SSD. Performance is shown as average request latency normalized to Fast-Only (i.e., PMEM device can accommodate all data).
Figure 13. Performance of Harmonia and baselines on an HSS with PMEM (intelpmem200, ) and high-end SSD (inteloptane, ). Performance is shown as average request latency normalized to Fast-Only. Lower is better.
We make four key observations. First, Harmonia outperforms all prior and extended techniques on the new HSS configuration. Harmonia provides an average performance improvement of 15.6%, 35.5%, 25.7%, 18.0%, 29.2%, 37.4%, 22.6% and 22.1% over CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM respectively, and achieves 80% of Oracle’s performance. Second, CDE is the best-performing prior approach in this HSS configuration because it prefers to place most data in the low-latency PMEM device. However, CDE’s frequent data evictions in the critical path lower its performance compared to Harmonia. CDE performs better in read-intensive workloads than in write-intensive workloads because read-intensive workloads consist of fewer writes and require fewer evictions. Third, in read-intensive workloads (i.e., RocksDB, MLPerf), Harmonia shows a higher performance improvement over prior approaches because it proactively migrates data to the PMEM device, which leads to more reads from the PMEM device. In write-intensive workloads, Harmonia’s benefits come from placing only performance-critical data in the PMEM device and migrating data across devices based on access patterns. Fourth, prior approaches without a proactive data-migration policy (e.g., Sibyl) have limited performance in read-intensive workloads because they rely on application updates to migrate data.
We discuss Harmonia’s applicability to emerging memory technologies in §7. We conclude that Harmonia provides significant performance benefits for HSS configurations with emerging memory technologies such as PMEM.

### v2

@@ 4.3. Coordination Between Harmonia’s Agents
We achieve coordination between the two agents through (1) their rewards, and (2) the effects of their actions on the HSS environment. The placement agent’s reward (See Equation 2) encourages effective utilization of the fast device, which is possible only if the migration agent proactively migrates cold (hot) data to the slow (fast) device. The data-migration agent receives a delayed reward (See Equation 3) that is based on access latencies of $n$ incoming I/O requests after migrating several candidate pages. In this way, each agent’s decisions influence the policy learned by the other agent.
Mitigating Conflicting Decisions. To prevent conflicting placement and migration decisions, we use two key techniques. First, the migration agent prioritizes the monitoring of pages with high access and migration intervals, and ignores recently placed or migrated pages (e.g., if the page’s access interval is 1, i.e., it was recently placed). Second, we add a small penalty based on access and migration intervals to the migration agent’s reward (See Equation 3) to discourage frequent page movements.
@@ 4.4. Harmonia: Design
Data-Placement Agent  1. An RL agent that places I/O request data in the best-fit storage device in the HSS 2 based on I/O request characteristics and HSS conditions. This agent consists of two threads. First, the data-placement training thread  3 is a background process in which the agent is trained. Second, the data-placement inference thread  4 is Harmonia’s only foreground thread. In this main decision thread, the data-placement agent decides the target storage device for incoming I/O request data on the critical path.
Data-Migration Agent 5. We implement the data-migration agent using two background threads to avoid interference with the I/O request handling on the critical path. First, the data-migration training thread  6 is used for training the agent. Second, in the data-migration inference thread  7, the agent performs three key tasks: (i) continuously monitors previously-placed pages in the HSS. (ii) determines migration candidates and their target devices. (iii) pushes the metadata of the migration candidates into the migration queue  8. A page is not a migration candidate if the target device determined by the migration agent is the same as its current device.
Migration Queue 8. A queue that stores the logical page address and the target device identifier (e.g., 0 for slow device) of each migration candidate identified by the migration agent. Based on empirical analysis (see §6.1), we set the migration queue size to 10 in our experiments to migrate pages in a timely manner during system idle times and avoid stale migration candidates.
Migration Thread 9. A background process that continuously monitors the migration queue and migrates pages across devices during system idle times. Harmonia samples device bandwidth at frequent intervals using real-time tools (e.g., iostat) to perform migrations. For high-intensity workloads with short system idle times, the migration thread checks if pages related to incoming read requests are already in the migration queue and issues low-priority writes (nvme, ; sata, ) to immediately schedule them for migration.
@@ 4.5. Design of Harmonia’s RL Agents
Inference Thread. Harmonia’s RL agents use the inference thread to make decisions while collecting experiences to use as training data. In this thread, the agent (1) observes the input features (state) a from the I/O request data and HSS, (2) performs inference using its inference network b, (3) selects the action c that maximizes d its long-term returns, and (4) collects the state, action and reward e as experiences f in an experience buffer g.
Training Thread. In this background thread, Harmonia uses collected experiences h from the experience buffer to train the training network i and optimize the placement or migration policy.
@@
Hyper-Parameter Tuning. We improve Harmonia’s accuracy by tuning the hyper-parameters (paine2020hyperparameter, ) of each agent using cross-validation (arlot2010survey, ) of different values. Harmonia’s agents need only one-time offline hyper-parameter tuning. Table 2 shows the hyper-parameter values chosen based on empirical analysis.
@@
Exploration vs. Exploitation. Harmonia’s RL agents start with no prior knowledge of the workload or HSS, and make random initial decisions (explore), and use their experiences (exploit) to gradually make optimal decisions. To balance exploration and exploitation, we use the $\epsilon$-greedy policy (tokic2011value, ): the predicted best action is selected with (1-$\epsilon$) probability, and a random action with $\epsilon$ probability.
Convergence of RL Agents in Harmonia. Figs 7(a) and  7(b) show the convergence of Harmonia’s data-placement and data-migration policies in terms of reduction in training loss for the first 10000 I/O requests in YCSB-B (see Table 4) on a performance-optimized HSS.
Figure 7. Convergence of Harmonia’s (a) Data-Placement and (b) Data-Migration RL agents shown in terms of reduction in training loss. We show only the first 10000 I/O requests.
We make two key observations: (1) The placement policy converges in less than 6000 I/O requests without prior training. (2) The migration policy takes longer to converge (around 9000 I/O requests) due to indirect and delayed reward to the migration agent. The convergence of Harmonia’s policies indicates that the agents learn to coordinate their actions and make optimal decisions.
@@ 4.6. Overhead Analysis
Harmonia has low inference and training latencies because of its (1) lightweight RL agent design, (2) separation of training and inference, and (3) background execution of multiple threads (e.g., data-migration inference thread). Harmonia’s two RL agents have identical networks. We perform training and inference on host CPU as the neural networks are lightweight, and their weights fit in the CPU on-chip caches in our evaluated system.
@@
Harmonia’s storage overhead includes (1) RL agents’ neural networks, (2) experience buffers, (3) migration queue, (4) migration agent’s reward, and (5) address mapping information. First, neural network weights are represented using a half-precision floating-point format. With 90 16-bit weights, each network requires $\sim$1.5 KiB of memory. The four networks (training and inference networks for two agents) take a negligible 6 KiB of memory, so we store their weights in the CPU caches. Second, to train each RL agent, we use an experience buffer located in the DRAM that consumes 100 KiB to store 1000 experiences. Hence, the total storage overhead of two experience buffers is 200 KiB. Overall, Harmonia’s RL agents require 206 KiB of memory. Third, the migration queue holds the logical block address (32 bits) and target device ID (4 bits) for each queue entry. The migration queue size is 10 in our experiments, requiring 50 bytes of DRAM. Fourth, we store the latencies of 50 incoming I/O requests temporarily to assign a delayed reward to the migration agent, which consumes  100B. Fifth, we maintain address mapping information for HSS devices (tsukada2021metadata, ). Harmonia requires 32 bits to store the state features for each page (see Table 1). This overhead is less than $\sim$0.8of the total storage capacity when using a 4-KiB data placement granularity.
@@ 5. Evaluation Methodology
We evaluate Harmonia on a real system with an HSS with two (dual-HSS), three (tri-HSS), and four (quad-HSS) devices. The dual-HSS configurations include: (1) performance-optimized HSS, which combines a high-end SSD (inteloptane, ) with a mid-range SSD (intels4510, ), and (2) cost-optimized HSS, which pairs the high-end SSD (inteloptane, ) with a low-end HDD (seagate, ). We implement a custom block device driver in Linux for Harmonia to interface with HSS devices. We implement Harmonia using the TF-Agents (TFAgents, ) library. Table 3 describes our real system and the HSS configurations.
@@ 1 TB, SATA 6Gb/s 7200 RPM
high-end (H) & middle-end (M) &
low-end SSD (LSSD) & low-end HDD (L)
Baselines. We compare Harmonia against four state-of-the-art heuristic- and machine-learning-based data-placement and migration techniques, (1) CDE (matsui2017design, ), (2) RNN-HSS (adapted from  (doudali2019kleio, )), (3) K-SVM (shetti2019machine, ), and (4) Sibyl (singh2022sibyl, ). We evaluate four extended approaches, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr and SAPM to explore the design space of holistic data-management techniques. We also compare against two ideal approaches: (1) Fast-Only, and (2) Oracle (meswani2015heterogeneous, ) (See § 3). Oracle serves as a reference to evaluate the accuracy of Harmonia’s decisions as it makes optimal decisions based on knowledge of future access patterns.
Workloads. We select seventeen data-intensive storage workloads from SYSTOR ’17 (lee2017understanding, ), RocksDB traces (yadgar2021ssd, ), Yahoo! Cloud Serving Benchmark (YCSB) suite (cooper2010benchmarking, ), and MLPerf Storage [^3] We generated the traces by running MLPerf applications on our evaluated system. (mlperfstoragebenchmarks, ) from real enterprise and datacenter environments. The average workload size is approximately 50000x the fast device capacity. We choose these workloads to represent diverse I/O access patterns with different read and write ratios, I/O request sizes, and inter-request times. Table 4 reports the characteristics of the chosen workloads. In our evaluation, each workload runs in a separate thread.
@@
Avg. Inter-
Request Time ($\mu$s)
@@
To evaluate Harmonia under real-world scenarios, we generate six multi-programmed workloads by running multiple workloads concurrently (see Table 5). We choose these multi-programmed workloads based on three key factors: (1) combination of read-intensive, write-intensive and mixed workloads, (2) number of concurrent workloads required for high I/O intensity (i.e., shorter inter-request times), and (3) combination of different request sizes. Each constituent workload is executed using a single thread. For example, mix6 runs on eight concurrent threads for the eight constituent workloads. Multi-programmed workloads typically exhibit higher I/O request intensity and varied access patterns.
@@ 6.1. Performance Analysis
Average I/O Request Latency and End-to-End Throughput. Figs. 8(a) and  8(b) show the performance of CDE, K-SVM, RNN-HSS, Sibyl, Sibyl+K-SVM, Sibyl+RNN-HSS, CDE+RL-Migr, SAPM, Harmonia and Oracle on performance- and cost-optimized HSS configurations. Figs. 9(a) and 9(b) show their end-to-end throughput on performance- and cost-optimized HSS, respectively.
Figure 8. Performance of Harmonia and baselines on performance-optimized (top) and cost-optimized (bottom) HSS configurations, shown as average request latency normalized to Fast-Only. Lower is better.
@@
We make four key observations. First, Harmonia consistently outperforms all baselines in both HSS configurations. In performance- (cost-) optimized HSS, Harmonia improves performance by 49.5% (31.7%) and end-to-end throughput by 49.4% (156.2%) over the best-performing prior approach, Sibyl, on average.
Second, Harmonia bridges the performance gap between Sibyl and Oracle by 64.2% (64.3%) and achieves 71.7% (62.2%) of Oracle’s throughput on average in performance- (cost-) optimized HSS.
Third, Harmonia shows higher performance and throughput benefits over baselines in read-intensive workloads (e.g., RocksDB, MLPerf) compared to write-intensive workloads. For example, Harmonia outperforms Sibyl by 57.5% (43.2%) in read-intensive workloads and 41.0% (39.9%) in write-intensive workloads. In read-intensive workloads, unlike Sibyl, Harmonia proactively prefetches frequently-read pages to the fast device to improve read performance even when there are no writes from the application, resulting in lower latencies and higher throughput. In write-intensive workloads, while both Harmonia and Sibyl use RL-based placement, Harmonia’s migration agent, in coordination with the placement agent, further improves performance by frequently migrating cold data from the fast device to accommodate new incoming requests.
Fourth, Harmonia shows higher performance and throughput gains over baselines (other than Sibyl) in cost-optimized than in performance-optimized HSS. This is because (1) data migration in cost-optimized HSS is more expensive than in performance-optimized HSS due to the large latency gap between the two devices, (2) prior approaches perform migration during I/O request handling, which adds significant latency overheads, and (3) Harmonia frees up space in the fast device with proactive migrations and performing migrations during system idle times. Sibyl’s RL agent places only performance-critical data in the fast device, reducing the number of data evictions on the critical path.
Tail Latency. Figs. 10(a) and 10(b) show the 99th and 99.99th percentile I/O latencies (tail latency) of Harmonia and baselines on performance-optimized HSS.
@@
We make two key observations. First, Harmonia significantly reduces tail latency compared to all baselines across the three workloads. Harmonia reduces the 99th (99.99th) percentile I/O latency by 32.3% (25.3%) on average over Sibyl. Harmonia has lower tail latency because, unlike prior techniques, Harmonia performs data migration during system idle times to avoid interference with ongoing I/O requests. Second, K-SVM, RNN-HSS, Sibyl+K-SVM and Sibyl+RNN-HSS show very high tail latencies because they perform a large volume of migration at fixed intervals, which adds a significant latency overhead to the ongoing I/O requests.
@@
Performance Sensitivity to Migration Queue Size. Fig.  11(a) shows the impact of migration queue size on HSS performance.
We make three key observations. First, optimal migration queue size is 10 for our evaluated workloads. This may vary slightly across workloads and HSS configurations. Second, low queue depths (i.e., ¡ 5) may result in many migration candidates being dropped, which leads to performance degradation. If the migration queue size is 0, Harmonia shows poor performance because it does not perform any data migration. Third, larger-sized queues can hold more migration candidates, but the oldest candidates may become stale (i.e., past migration decisions may not correlate with current access patterns and HSS conditions), which adversely impacts HSS performance.
Performance Sensitivity to Migration Agent’s Reward. Fig.  11(b) shows the performance impact of the number of I/O requests considered for the migration agent’s delayed reward (See Equation 3).
@@
We make three key observations. First, the optimal number of requests is 50 in our evaluated configuration, but may vary slightly across workloads and HSS configurations, and depends on factors including (1) inter-request times of incoming I/O requests after migration, (2) workload access patterns (e.g., access frequency and reuse distance of a page), and (3) system conditions (e.g., device capacity utilization). Second, considering fewer requests (e.g., ¡ 10) may not provide the right reward to the agent because the effect of migrations is not seen immediately in the placement latencies of these few requests. Third, if we consider more I/O requests (e.g., 250), we see lower performance because of two reasons: (i) the impact of data migration is diluted if we consider a large number of incoming requests, and (ii) I/O requests that arrive too far into the future after data migration may not see a direct impact.
Performance on HSS with Persistent Memory (PMEM) Devices. To demonstrate the Harmonia’s extensibility to emerging persistent memory technologies, we evaluate Harmonia and baselines on an HSS with a persistent memory device (intelpmem200, ) and a high-end SSD (inteloptane, ) (see Table 3). We emulate the PMEM device using RAMDisk because these emerging devices are not easily available. Fig. 12 shows the performance of Harmonia and baselines on HSS with PMEM. Performance is shown as average request latency normalized to Fast-Only (i.e., PMEM device can store all data).
Figure 12. Performance of Harmonia and baselines on an HSS with PMEM (intelpmem200, ) and high-end SSD (inteloptane, ) shown as average request latency normalized to Fast-Only. Lower is better.
We make three key observations. First, Harmonia outperforms the best-performing baseline, CDE, by 15.6% and achieves 80% of Oracle’s performance. Second, CDE outperforms other baselines in this HSS configuration as it places most data in the low-latency PMEM device. However, CDE’s frequent data evictions on the critical path lower its performance compared to Harmonia. CDE performs better in read-intensive workloads because they consist of fewer writes and require fewer evictions. Third, in read-intensive workloads (i.e., RocksDB, MLPerf), Harmonia shows a higher performance improvement as it proactively prefetches data to the PMEM device, which leads to more reads from the PMEM device. We discuss Harmonia’s applicability to emerging memory technologies in §7.